Two numbers have a sum of 36, a difference of 24, and a product of 180. Dividing the greater one by the lesser one gives a quotient of 5. What are the numbers.
I keep thinking it has to do with the numbers 36 & 5 but they don't fit all the situations.

Answer by mananth(16946)

Two numbers have a sum of 36, a difference of 24, and a product of 180. Dividing the greater one by the lesser one gives a quotient of 5. What are the numbers.

let the numbers be x & y
x+y=36
x-y=24
x+y+x-y= 60
2x=60
x=30
the other number = 6